# Fermion and Boson Pairs in Beamsplitters and MZIs

### v1

@@ Abstract
In this short Topical Review, we look at something typically considered trivial, but not given formally elsewhere —the behaviour of first multiple fermions then multiple bosons at a beamsplitter, and, extending from this, the behaviour of multiple fermions and multiple bosons in Mach-Zehnder interferometers (MZIs). We hope that by showing how to go from mathematically-simple but unintuitive quantum field theory to a phenomenological description, this Review will help researchers and students in the field build a stronger intuition for the behaviour of quantum particles.
@@ I Introduction
It is regularly said that quantum field theory is just linear algebra, at least for non-continuous degrees of freedom (be they finite, e.g., polarisation or spin, or nominally infinite e.g., photon number or orbital angular momentum). It is therefore almost as often assumed that it is simple to work out the behaviour of these systems - after all, most people learn linear algebra in high school. However, given how surprising certain key results seem (e.g., the Hong-Ou-Mandel effect or “HOM dip” [1]) [^1] Combined with the hours of confusion at the recent Photon 2024, which resulted from asking photonics theorists and experimentalists what would happen if you put two indistinguishable photons into the same port of a Mach-Zehnder interferometer, then changed the phase between the two arms., these effects are possibly not as obvious as often claimed.
Therefore, we decided it would be best to elucidate the behaviour of two of the scenarios which raise the most confusion—what happens if you put a fermion into a beamsplitter [2, 3], acting on it the same way a beamsplitter acts on a photon; and what happens if you put either two fermions, or two bosons, of some level of indistinguishability, into a Mach-Zehnder interferometer. We aim for this treatment to be as thorough as possible, also considering the effects of adding phases between different paths, or distinguishability between the different bosons, in order to help readers develop an intuition for these scenarios, and for boson and fermion interferometry in general, in the same vein as others who have started this quest (e.g., Loudon, in [4]). The underpinning theory behind everything in this review can be found in standard quantum optics textbooks, such as [5, 6, 7], yet the effects described link to one of the most peculiar aspects of quantum mechanics—quantum interference.
@@ II.1 Two Indistinguishable Particles arriving Simultaneously
For the simplest case (a 50:50) beamsplitter, this gives
@@
$\left|\psi_{f}\right\rangle=\frac{1}{2}\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)\left|0_{c},0_{d}\right\rangle$ (4)
@@
$\begin{split}\left|\psi_{f}\right\rangle_{\textrm{boson}}&=\frac{1}{2}\left((\hat{c}^{\dagger})^{2}-\hat{c}^{\dagger}\hat{d}^{\dagger}+\hat{d}^{\dagger}\hat{c}^{\dagger}-(\hat{d}^{\dagger})^{2}\right)\left|0_{c},0_{d}\right\rangle\\ &=\frac{1}{2}\left((\hat{c}^{\dagger})^{2}-(\hat{d}^{\dagger})^{2}\right)\left|0_{c},0_{d}\right\rangle\\ &=\frac{1}{\sqrt{2}}\left(\left|2_{c},0_{d}\right\rangle-\left|0_{c},2_{d}\right\rangle\right)\end{split}$ (6)
@@
$\begin{split}\left|\psi_{f}\right\rangle_{\textrm{fermion}}&=\frac{1}{2}\left((\hat{c}^{\dagger})^{2}-\hat{c}^{\dagger}\hat{d}^{\dagger}+\hat{d}^{\dagger}\hat{c}^{\dagger}-(\hat{d}^{\dagger})^{2}\right)\left|0_{c},0_{d}\right\rangle\\ &=\frac{1}{2}\left(-\hat{c}^{\dagger}\hat{d}^{\dagger}-\hat{c}^{\dagger}\hat{d}^{\dagger}\right)\left|0_{c},0_{d}\right\rangle\\ &=\frac{-2}{2}\hat{c}^{\dagger}\hat{d}^{\dagger}\left|0_{c},0_{d}\right\rangle=-\left|1_{c},1_{d}\right\rangle\end{split}$ (11)
@@ II.2 Adding Distinguishability between the Particles
Following [8], let’s quantify distinguishability between the two particles with parameter $r_{i,j}$, a real number, which is one when the particles are indistinguishable, and zero when they are perfectly distinguishable.
What does this parameter do? Well, we can view it as how similar the creation operators in a given mode are for our two particles; i.e., for our second particle, its creation operator in some mode will be a weighted sum of the same creation operator for that mode as for particle 1, and an orthogonal creation operator:
@@
$\begin{split}\left|\psi\right\rangle&=\hat{a}^{\dagger}_{1}\hat{b}^{\dagger}_{2}\left|0,0\right\rangle=\frac{1}{2}\left(\hat{c}^{\dagger}_{1}+\hat{d}^{\dagger}_{1}\right)\left(\hat{c}^{\dagger}_{2}-\hat{d}^{\dagger}_{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{2}\left(\hat{c}^{\dagger}_{1}+\hat{d}^{\dagger}_{1}\right)\left(\sqrt{r_{1,2}}\hat{c}^{\dagger}_{1}+\sqrt{1-r_{1,2}}\hat{c}^{\dagger}_{1\perp}-\sqrt{r_{1,2}}\hat{d}^{\dagger}_{1}-\sqrt{1-r_{1,2}}\hat{d}^{\dagger}_{1\perp}\right)\left|0,0\right\rangle\\ &=\frac{1}{2}\left(\sqrt{r_{1,2}}\left((\hat{c}^{\dagger}_{1})^{2}-\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{1}+\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{1}-(\hat{d}^{\dagger}_{1})^{2}\right)+\sqrt{1-r_{1,2}}\left(\hat{c}^{\dagger}_{1}\hat{c}^{\dagger}_{1\perp}-\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{1\perp}+\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{1\perp}-\hat{d}^{\dagger}_{1}\hat{d}^{\dagger}_{1\perp}\right)\right)\left|0,0\right\rangle\end{split}$ (19)
@@ III Mach-Zehnder Interferometers
Figure 2: A Mach-Zehnder interferometer, formed of two 50:50 beamsplitters (and two mirrors). Note here that while mode $a$ connects to mode $c$ by transmission in the first beamsplitter, in the second beamsplitter they connect by reflection. In practice, this doesn’t imply the beamsplitters are different, but reflects the effects of phases caused by the two mirrors.
@@ III.1 Indistinguishable Particles
$\left|\psi\right\rangle=\frac{1}{2}\left((\hat{c}^{\dagger})^{2}-\hat{c}^{\dagger}\hat{d}^{\dagger}+\hat{d}^{\dagger}\hat{c}^{\dagger}-(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle$ (20)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2}\left((\hat{a}^{\dagger})^{2}-\hat{a}^{\dagger}\hat{b}^{\dagger}+\hat{b}^{\dagger}\hat{a}^{\dagger}-(\hat{b}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4}\left(\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)^{2}-\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)+\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)-\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)^{2}\right)\left|0,0\right\rangle\\ &=\hat{c}^{\dagger}\hat{d}^{\dagger}\left|0,0\right\rangle=\left|1,1\right\rangle\end{split}$ (21)
@@
From a mathematical perspective this makes sense, given the matrix we gave above for a 50:50 phaseless beamsplitter, when multiplied by itself, gives us the identity matrix—but physically this still seems peculiar. Regardless if the particles bunch or antibunch, for this interferometer, if two indistinguishable particles go in via different ports, they always come out via different ports. This leads us to ask: is this affected by the phase between different components in the interferometer? (Note, as we mention above, this phase could be caused by there being some phase in the beamsplitter; or, it could be caused by adding a phase to one of the two paths, e.g., using a quarter-wave plate.)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{1}{2}\left((\hat{a}^{\dagger})^{2}-e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}-e^{2i\phi}(\hat{b}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4}\left(\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)^{2}-e^{i\phi}\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)+e^{i\phi}\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)-e^{2i\phi}\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+(1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1-e^{i\phi})^{2}\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\end{split}$ (22)
As expected, at $\phi=0$, this gives us the behaviour we see above. When $\phi=\pi$, we see the same behaviour—but this is from the $\hat{d}^{\dagger}\hat{c}^{\dagger}$ component, rather than the $\hat{c}^{\dagger}\hat{d}^{\dagger}$. At phases other than these, we get a superposition of the terms, which require us to consider whether we are putting bosons or fermions into the interferometer.
@@
$\begin{split}\left|\psi\right\rangle_{\textrm{bosons}}&=\frac{1}{4}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+(1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+2(1+e^{2i\phi})\hat{c}^{\dagger}\hat{d}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{(1-e^{2i\phi})}{2\sqrt{2}}\left(\left|2,0\right\rangle+\left|0,2\right\rangle\right)+\frac{(1+e^{2i\phi})}{2}\left|1,1\right\rangle\end{split}$ (23)
@@
$\begin{split}\left|\psi\right\rangle_{\textrm{fermions}}&=\frac{1}{4}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+(1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1-e^{i\phi})^{2}\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4}\left((1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}-(1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}\right)\left|0,0\right\rangle\\ &=e^{i\phi}\left|1,1\right\rangle\end{split}$ (24)

### v2

@@ Abstract
In this short Topical Review, we look at something typically considered trivial, but not given formally elsewhere —the behaviour of first multiple fermions, then multiple bosons, at a beamsplitter. Extending from this, we then describe the behaviour of multiple fermions and multiple bosons in Mach-Zehnder interferometers (MZIs). We hope that by showing how to go from mathematically-simple but unintuitive quantum field theory to a phenomenological description, this Review will help both researchers and students build a stronger intuition for the behaviour of quantum particles.
@@ I Introduction
It is often said that quantum field theory is just linear algebra. This is true at least for non-continuous degrees of freedom, be they finite, e.g., polarisation or spin, or nominally infinite e.g., photon number or orbital angular momentum. It is therefore almost as often assumed that it is simple to work out the behaviour of these systems—after all, most people learn linear algebra in high school. However, given how surprising certain key results seem (e.g., the Hong-Ou-Mandel effect or “HOM dip” [1]) [^1] Combined with the hours of confusion at the recent Photon 2024, which resulted from asking photonics theorists and experimentalists what would happen if you put two indistinguishable photons into the same port of a Mach-Zehnder interferometer, then changed the phase between the two arms., these effects are possibly not as obvious as claimed.
Therefore, we decided it would be best to consider the behaviour of two of the scenarios which raise the most confusion—what happens if you put a fermion into a beamsplitter [2, 3], acting on it the same way a beamsplitter acts on a photon; and what happens if you put either two fermions, or two bosons, of some level of indistinguishability, into a Mach-Zehnder interferometer. We aim for this treatment to be as thorough as possible, meaning we also considering the effects of adding phases between different paths, and distinguishability between the different bosons. We do this in order to help readers develop an physical intuition for these scenarios, and for boson and fermion interferometry in general. This paper is therefore in the same vein as earlier works which aim to help the reader develop a physical intuition for quantum-field-theoretic effects (e.g., Loudon, in [4]). Even though the underpinning theory behind everything in this review can be found in standard quantum optics textbooks, such as [5, 6, 7], the effects described link to one of the most peculiar aspects of quantum mechanics—quantum interference.
@@ II.1 Two Indistinguishable Particles arriving Simultaneously
where $\theta$ parametrises the transmissivity/reflectivity of the beamsplitter: the beamsplitter is perfectly transmissive when $\theta=0$, perfectly reflective when $\theta=\pi$, and equally transmissive and reflective in the situations we consider, where $\theta=\pi/2$ (see e.g., Chapter 5.1 of [7] for a more general treatment of beamsplitters). In these situations, this gives
@@
which we can rearrange to give
$\begin{split}\hat{a}^{\dagger}=\frac{1}{\sqrt{2}}(\hat{c}^{\dagger}+\hat{d}^{\dagger}),\;\hat{b}^{\dagger}=\frac{-1}{\sqrt{2}}(\hat{c}^{\dagger}-\hat{d}^{\dagger})\end{split}$ (3)
@@
$\left|\psi_{f}\right\rangle=\frac{-1}{2}\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)\left|0_{c},0_{d}\right\rangle$ (5)
@@
$\begin{split}\left|\psi_{f}\right\rangle_{\textrm{boson}}&=\frac{-1}{2}\left((\hat{c}^{\dagger})^{2}-\hat{c}^{\dagger}\hat{d}^{\dagger}+\hat{d}^{\dagger}\hat{c}^{\dagger}-(\hat{d}^{\dagger})^{2}\right)\left|0_{c},0_{d}\right\rangle\\ &=\frac{-1}{2}\left((\hat{c}^{\dagger})^{2}-(\hat{d}^{\dagger})^{2}\right)\left|0_{c},0_{d}\right\rangle\\ &=\frac{1}{\sqrt{2}}\left(\left|0_{c},2_{d}\right\rangle-\left|2_{c},0_{d}\right\rangle\right)\end{split}$ (7)
@@
$\begin{split}\left|\psi_{f}\right\rangle_{\textrm{fermion}}&=\frac{-1}{2}\left((\hat{c}^{\dagger})^{2}-\hat{c}^{\dagger}\hat{d}^{\dagger}+\hat{d}^{\dagger}\hat{c}^{\dagger}-(\hat{d}^{\dagger})^{2}\right)\left|0_{c},0_{d}\right\rangle\\ &=\frac{-1}{2}\left(-\hat{c}^{\dagger}\hat{d}^{\dagger}-\hat{c}^{\dagger}\hat{d}^{\dagger}\right)\left|0_{c},0_{d}\right\rangle\\ &=\frac{2}{2}\hat{c}^{\dagger}\hat{d}^{\dagger}\left|0_{c},0_{d}\right\rangle=\left|1_{c},1_{d}\right\rangle\end{split}$ (12)
@@ II.2 Adding Distinguishability between the Particles
Following [8], let’s quantify distinguishability between the two particles with parameter $r_{i,j}$, a real number, which is one when the particles are indistinguishable, and zero when they are perfectly distinguishable. In a real experiment, this partial distinguishability could arise in many different ways: it could come from differences in polarisation between the two particles (e.g., if particle 1 was $H$-polarised, this could represent particle 2 being in some combination (be it superposition or mixture) of $H$- and $V$-polarised); alternatively, it could be a way of accounting for the particles not arriving at the beamsplitter exactly simultaneously, or slight near- or far-field misalignment, or one of many other ways the particles could be made slightly distinguishable. All that matters though here, for our purposes, is that they are slightly distinguishable, in a way that can be quantified through $r_{1,2}$.
What does this parameter $r_{1,2}$ do? Well, we can view it as how similar the creation operators in a given mode are for our two particles; i.e., for our second particle, its creation operator in some mode will be a weighted sum of the same creation operator for that mode as for particle 1, and an orthogonal creation operator:
@@
$\begin{split}\left|\psi\right\rangle&=\hat{a}^{\dagger}_{1}\hat{b}^{\dagger}_{2}\left|0,0\right\rangle=\frac{-1}{2}\left(\hat{c}^{\dagger}_{1}+\hat{d}^{\dagger}_{1}\right)\left(\hat{c}^{\dagger}_{2}-\hat{d}^{\dagger}_{2}\right)\left|0,0\right\rangle\\ &=\frac{-1}{2}\left(\hat{c}^{\dagger}_{1}+\hat{d}^{\dagger}_{1}\right)\left(\sqrt{r_{1,2}}\hat{c}^{\dagger}_{1}+\sqrt{1-r_{1,2}}\hat{c}^{\dagger}_{1\perp}-\sqrt{r_{1,2}}\hat{d}^{\dagger}_{1}-\sqrt{1-r_{1,2}}\hat{d}^{\dagger}_{1\perp}\right)\left|0,0\right\rangle\\ &=\frac{-1}{2}\left(\sqrt{r_{1,2}}\left((\hat{c}^{\dagger}_{1})^{2}-\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{1}+\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{1}-(\hat{d}^{\dagger}_{1})^{2}\right)+\sqrt{1-r_{1,2}}\left(\hat{c}^{\dagger}_{1}\hat{c}^{\dagger}_{1\perp}-\hat{c}^{\dagger}_{1}\hat{d}^{\dagger}_{1\perp}+\hat{d}^{\dagger}_{1}\hat{c}^{\dagger}_{1\perp}-\hat{d}^{\dagger}_{1}\hat{d}^{\dagger}_{1\perp}\right)\right)\left|0,0\right\rangle\end{split}$ (20)
@@ III Mach-Zehnder Interferometers
Figure 2: A Mach-Zehnder interferometer, formed of two 50:50 beamsplitters (and two mirrors).
@@ III.1 Indistinguishable Particles
$\left|\psi\right\rangle=\frac{-1}{2}\left((\hat{c}^{\dagger})^{2}-\hat{c}^{\dagger}\hat{d}^{\dagger}+\hat{d}^{\dagger}\hat{c}^{\dagger}-(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle$ (21)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{-1}{2}\left((\hat{a}^{\dagger})^{2}-\hat{a}^{\dagger}\hat{b}^{\dagger}+\hat{b}^{\dagger}\hat{a}^{\dagger}-(\hat{b}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{1}{4}\left(-\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)^{2}-\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)+\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)+\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)^{2}\right)\left|0,0\right\rangle\\ &=-\hat{d}^{\dagger}\hat{c}^{\dagger}\left|0,0\right\rangle=-\left|1,1\right\rangle\end{split}$ (22)
@@
From a mathematical perspective this makes sense, given the matrix we gave above for a 50:50 phaseless beamsplitter, when multiplied by itself, gives us $-i$ times the Pauli-$y$ matrix (so together we would expect from them a flip in modes, with a negative phase placed on one of the modes). However, physically, this still seems peculiar. Regardless if the particles bunch or antibunch, for this interferometer, if two indistinguishable particles go in via different ports, they always come out via different ports. This leads us to ask: is this affected by the phase between different components in the interferometer? (Note, as we mention above, this phase could be caused by there being some phase in the beamsplitter; or, it could be caused by adding a phase to one of the two paths, e.g., using a quarter-wave plate.)
@@
$\begin{split}\left|\psi\right\rangle&=\frac{-1}{2}\left((\hat{a}^{\dagger})^{2}-e^{i\phi}\hat{a}^{\dagger}\hat{b}^{\dagger}+e^{i\phi}\hat{b}^{\dagger}\hat{a}^{\dagger}-e^{2i\phi}(\hat{b}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{-1}{4}\left(\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)^{2}+e^{i\phi}\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)-e^{i\phi}\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)\left(\hat{c}^{\dagger}+\hat{d}^{\dagger}\right)-e^{2i\phi}\left(\hat{c}^{\dagger}-\hat{d}^{\dagger}\right)^{2}\right)\left|0,0\right\rangle\\ &=\frac{-1}{4}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+(1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1+e^{i\phi})^{2}\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\end{split}$ (23)
As expected, at $\phi=0$, this gives us the behaviour we see above. When $\phi=\pi$, we see the same behaviour—but this is from the $\hat{c}^{\dagger}\hat{d}^{\dagger}$ component, rather than the $\hat{d}^{\dagger}\hat{c}^{\dagger}$. At phases other than these, we get a superposition of the terms, which require us to consider whether we are putting bosons or fermions into the interferometer.
@@
$\begin{split}\left|\psi\right\rangle_{\textrm{bosons}}&=\frac{-1}{4}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+(1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{-1}{4}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+2(1+e^{2i\phi})\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{(1-e^{2i\phi})}{2\sqrt{2}}\left(\left|2,0\right\rangle+\left|0,2\right\rangle\right)+\frac{(1+e^{2i\phi})}{2}\left|1,1\right\rangle\end{split}$ (24)
@@
$\begin{split}\left|\psi\right\rangle_{\textrm{fermions}}&=\frac{-1}{4}\left((1-e^{2i\phi})(\hat{c}^{\dagger})^{2}+(1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}+(1+e^{i\phi})^{2}\hat{d}^{\dagger}\hat{c}^{\dagger}+(1-e^{2i\phi})(\hat{d}^{\dagger})^{2}\right)\left|0,0\right\rangle\\ &=\frac{-1}{4}\left((1-e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}-(1+e^{i\phi})^{2}\hat{c}^{\dagger}\hat{d}^{\dagger}\right)\left|0,0\right\rangle\\ &=-e^{i\phi}\left|1,1\right\rangle\end{split}$ (25)